OpenStudy (anonymous):

Find exact value: tan(π/6 + π/4) how to plug into Sum and difference formula

OpenStudy (anonymous):

\[\tan(\frac{ \pi }{ 6} + \frac{ \pi }{ 4 }) \]

OpenStudy (anonymous):

using formula \[\tan \left( a + b \right)=\left(\begin{matrix}\frac{ tanA +tanB }{ 1- tanA \times tanB} \\ \end{matrix}\right)\]

OpenStudy (anonymous):

You have all the information you need, just find the tan(pi/6) and the tan(pi/4) then plug in your values into the formula.

OpenStudy (anonymous):

i dont know what value is tan ... I know how to do sin / cos but not tan..

OpenStudy (anonymous):

*tan is

OpenStudy (anonymous):

Well, just take your value of sine, and divide it by cosine, because that is the definition of tangent.

OpenStudy (anonymous):

\[\tan \frac{ \pi }{ 6 }=\frac{ \sin \frac{ \pi }{ 6 } }{ \cos \frac{ \pi }{ 6 } }\]

OpenStudy (anonymous):

and do the same for tan(pi/4)

OpenStudy (anonymous):

If you have your unit circle committed to memory, then all you have to do is take your y @ 30 degrees, and divide by your x @ 30* to find your tan(pi/6).

OpenStudy (anonymous):

sorry I had to take a break for a min. thank you!

OpenStudy (anonymous):

No Problem. Anytime.

OpenStudy (anonymous):

:( ... I still dont understand how to set this up. I answer key say answer is \[\frac{ \sqrt{3}+ 1 }{ \sqrt{3} - 1}\]

OpenStudy (anonymous):

*the answer key states that correct answer is:

OpenStudy (anonymous):

ok, let's compute each expression piece by piece.

OpenStudy (anonymous):

Takes me forever to use equation tool

OpenStudy (anonymous):

tan(pi/6) = (1/2) / (sqrt3 / 2)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so the 2's cancel, leaving you 1 over rad 3.

OpenStudy (anonymous):

tan(pi/4)= (sqrt2)/(sqrt2)

OpenStudy (anonymous):

which simplifies to 1.

OpenStudy (anonymous):

So, this is what we have: \[\frac{ 1-\frac{ 1 }{ \sqrt{3} } }{ 1-\frac{ 1 }{ \sqrt{3} } } \]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

I mean + on top.

OpenStudy (anonymous):

1+1/rad 3

OpenStudy (anonymous):

ok, lemme start this again so as not to confuse you.

OpenStudy (anonymous):

not that hard at all. haha. Just looks intimidating

OpenStudy (anonymous):

ok, so did you understand how the final result was obtained or do you need to see the steps?

OpenStudy (anonymous):

I understand, cheers!

OpenStudy (anonymous):

great! Glad I helped...somewhat...lol